OpenStudy (anonymous):

\documentclass[11pt,oneside,letter]{article} \usepackage{amsmath} \usepackage{mathtools} \begin{document} \begin{align*} \frac{\delta y}{\delta x} = \int{x^2 \sin(x) dx} \end{align*} \end{document}

OpenStudy (anonymous):

\[\frac{\delta y}{\delta x} = \int{x^2 \sin(x)}dx\]

OpenStudy (anonymous):

\( \text{The answer to life, the universe and everything is }42.\)

OpenStudy (anonymous):

\(\text{\bfseries This sentence is bold}\)

OpenStudy (anonymous):

\(\color{blue}{\text{This sentence is blue.}}\)

OpenStudy (anonymous):

\(\LaTeX{} \)

OpenStudy (experimentx):

??